# Aligning Perception, Reasoning, Modeling and Interaction: A Survey on Physical AI

### v3

@@ F.2 Navigation
Object-Goal Navigation. Object-Goal Navigation (ObjectNav) tasks an agent to interact with its physical world by exploring an unknown space to find a target object, requiring it to reason about potential object locations based on visual cues. It’s best represented by the Habitat-Matterport3D (MP3D) [239] benchmark, and benchmarks like AI2-THOR [240], Gibson [241], RoboTHOR [242], HM3D-OVON [243] belong to this task.
Vision-and-Language Navigation. Vision-and-Language Navigation (VLN) challenges an agent to reason about the connection between language and the physical world, interacting with its environment by translating natural language commands into a specific navigation path. For example, R2R [244, 239], RxR [245], REVERIE [246] and TOUCHDOWN [247] are benchmarks for this task.
Dialog-based Navigation. Dialog-based Navigation allows an agent to interact with a user and its physical world simultaneously, reasoning about ambiguous instructions through conversation to successfully navigate its environment. Related to it are RobotSlang [248], R2H [249], CVDN [250], UNMuTe [251] and so on.
Evolution of Navigation Methods. Navigation tasks have seen a significant methodological evolution, moving from early, specialized systems to models powered by large language models (LLMs). Early non-LLM-based methods were typically trained on domain-specific data to create a direct mapping from sensory inputs to navigation actions [252]. This reliance on implicit representations often led to models that learned spurious correlations between inputs and outputs, limiting their generalization to unseen environments and making their decisions difficult to interpret [253].
The advent of LLMs and VLMs has introduced a new paradigm. Pre-trained on vast, diverse datasets, these models bring a wealth of real-world commonsense and physical knowledge that dramatically improves navigation precision. This shift has enabled explicit reasoning, such as CoT [57], which not only enhances decision accuracy but also provides greater interpretability [253]. This has led to a new wave of methods which are shown in Table III and Figure 8.
Figure 8: Overview of NavCoT (Figure used courtesy of [254]).
TABLE III: Representative methods for navigation tasks: Non-LLM vs. LLM-based
@@
ObjectNav
GOSE [255]
RIM [256]
LOAT [257]
LGR [258]
CL-CoTNav [259]
ASCENT [260]
VLN
BabyWalk [261]
ADAPT [262]
HAMT [263]
EnvDrop [264]
DUET [265]
NavGPT [266]
NavCoT [254]
VELMA [267]
NaVILA [268]
Dialog-based
CMN [269]
GVDN [270]
LLM-
ContextBridge [271]
FLAME [272]
However, these LLM-based approaches are not without their issues. They are susceptible to hallucinations, where incorrect physical reasoning can lead to failures, such as imagining a non-existent path through an obstacle [273]. This fragility is further demonstrated by research showing that minor changes to a prompt can reduce a robot’s task success rate by nearly 20% [274]. A fundamental limitation is the representational bottleneck caused by converting rich physical world data into a token format, which can result in a significant loss of detail and the inability to reason about low-level preconditions for actions [256, 255]. This is addressed by integrating world models, which can mentally simulate action outcomes, providing a crucial, grounded, and predictive layer that LLMs inherently lack [275]. This paradigm is also essential for bridging the "sim-to-real" gap, as a world model can learn a robust predictive representation of dynamics in simulation and transfer this knowledge to a real-world agent, as demonstrated by the TWIST [276] framework’s teacher-student distillation approach.
@@ F.3 Autonomous Driving
Autonomous driving refers to a system’s capability to perform part or all of the dynamic driving task (DDT) on a sustained basis without direct human intervention. The DDT includes all of the real-time operational and tactical functions required to operate a vehicle, such as steering, acceleration, and braking. To achieve this, an autonomous vehicle’s closed-loop software system integrates a series of core functions: perception, prediction, planning, and control [277]. Perception involves collecting and processing real-time data from various sensors like cameras and LiDAR to detect objects and road conditions [278]. Prediction forecasts the behavior of other road users [279]. Planning generates a safe and efficient path for the vehicle to follow. Control translates the planned path into physical commands for the vehicle’s actuators [280]. These tasks are implemented through several architectural paradigms [281]. Current autonomous driving technology methods mainly include: Rule-based Approaches, Learning-based Approaches, World Models and Generative Approaches, Hybrid Approaches, etc.
@@
Rule-based Approaches. Rule-based approaches rely on deterministic algorithms and explicit domain knowledge, encoding physics and traffic rules in interpretable models. Classical methods include the Intelligent Driver Model (IDM) [282] for car-following, RRT/RRT* [283] for sampling-based trajectory generation, Model Predictive Control (MPC) [284] for optimization-based planning, and potential-field navigation [285]. They are transparent and safe but struggle in dense and uncertain traffic. Benchmarks include CommonRoad [286], nuPlan [287], nuScenes [288], Waymo Open Dataset (WOD) [289], KITTI [290], Cityscapes [291], Argoverse [292], Lyft Level 5 [293], ApolloScape [294], BDD100K [295], CARLA [296], and AirSim [297].
@@
Learning-based Approaches. Learning-based approaches replace rules with neural policies trained from data. Imitation Learning (IL) follows expert trajectories (e.g., ChauffeurNet [298]), Inverse Reinforcement Learning (IRL) recovers reward functions from demonstrations [299], and Reinforcement Learning (RL) optimizes policies in simulators [300]. AlphaDrive [21] and AutoDrive-R2 [22] extend RL-based training with physics-informed rewards, reinforcing the role of simulation and safety-oriented benchmarks in this category. More recent methods use transformer-based forecasting (e.g., LaneGCN [301], Wayformer [302]).
@@
Generative Approaches. World models learn latent spatio-temporal representations of the driving environment, enabling rollouts of future states [1]. The Dreamer family (DriveDreamer [303] in Figure 9, DriveDreamer4D [304], ReconDreamer [305]) captures compact latent dynamics for long-horizon simulation. Diffusion-based models include DrivingDiffusion [306], BEVControl [307], and Vista [308]. Transformer-based models such as GAIA-1/2 [184], HoloDrive [309], DrivingWorld [310] and DrivingGPT [311] support multimodal controllable scene generation. Occupancy-based models predict 4D voxelized space, such as OccWorld [312], OccLLaMA [313], RenderWorld [314], and the newly proposed Drive-OccWorld [315] and DOME [316], which combine occupancy forecasting with planning via diffusion. Point-cloud generators like LiDARGen [317], Copilot4D [318], UltraLiDAR [319], and the unified HERMES [320] framework produce future LiDAR sweeps or merged 3D representations. Very recent works include DreamDrive [321] for 4D generative modeling and STAG-1 [322]for realistic video-driven 4D scene simulation.
Benchmarks in this domain include prediction and occupancy forecasting benchmarks (Occ3D-nuScenes/Waymo [323], OpenScene [324], KITTI-360 [325], Argoverse 2 [326], Precog [327]), simulation benchmarks (CarlaSC [328], Bench2Drive [329]), and end-to-end planning benchmarks (nuPlan [287], S2R-Bench [330]). These collectively test world models across scene generation, motion forecasting, and driving policy integration.
Figure 9: Overall framework of DriveDreamer (Figure used courtesy of [303]).
@@
Hybrid Approaches. Hybrid approaches combine rule-based reliability with learning-based adaptability and cognitive reasoning. Neural waypoint predictors refined by MPC or rule-based controllers augmented with learned forecasting exemplify this paradigm [331]. Cognitive reasoning extensions include DriveCoT [332], which integrates Chain-of-Thought supervision into driving tasks, and PRIMEDrive-CoT [333], which introduces uncertainty-aware reasoning. PlanAgent separates high-level planning from execution, while LeapVAD [334] adopts a dual-process architecture inspired by human cognition. DriveLMM-o1 [335] and Reason2Drive [336] further contribute cognition-augmented datasets for training and evaluating reasoning.
@@
The current challenges in autonomous driving are fundamentally tied to the "long tail" problem of rare, high-impact edge cases that are difficult to encounter and address through traditional real-world data collection [337, 338, 339]. World models solve this by creating a learned, internal simulation of the environment, which functions as a "computational snow globe" where the AI can "mentally rehearse" actions [340, 341]. By generating and training in an unbounded number of synthetic, safety-critical scenarios, these models allow the autonomous system to anticipate and mitigate hazardous situations before they unfold in the physical world, ultimately shifting the paradigm from reactive to predictive autonomy and improving real-world safety [342].
@@ H Conclusion
This survey presents a comprehensive task-oriented taxonomy of AI systems for physics reasoning, organizing the landscape into four fundamental capabilities: physical perception, physics reasoning, world modeling, and embodied interaction. While our analysis reveals significant progress in specialized domains, most existing systems still rely on statistical pattern recognition rather than genuine physical understanding, with the most promising developments emerging from hybrid approaches that integrate physics-grounded architectures, physics-informed training methodologies, and explicit symbolic reasoning. The field stands at a critical juncture where recent breakthroughs in large language models, Graph Neural Networks, and multimodal systems offer unprecedented opportunities to develop unified frameworks that seamlessly bridge theoretical physics reasoning with embodied physical interaction. As AI systems increasingly operate in complex physical environments, the integration of robust physical reasoning capabilities becomes essential for safety, reliability, and interpretability, requiring continued collaboration between AI researchers, physicists, and domain experts to realize truly physical intelligent systems that comprehend and interact with physical reality.

### v4

@@ F.2 Navigation
Object-Goal Navigation. Object-Goal Navigation (ObjectNav) tasks an agent to interact with its physical world by exploring an unknown space to find a target object, requiring it to reason about potential object locations based on visual cues. It’s best represented by the Habitat-Matterport3D (MP3D) [Matterport3D] benchmark, and benchmarks like AI2-THOR [AI2-THOR], Gibson [Gibson], RoboTHOR [RoboTHOR], HM3D-OVON [HM3D-OVON] belong to this task.
Vision-and-Language Navigation. Vision-and-Language Navigation (VLN) challenges an agent to reason about the connection between language and the physical world, interacting with its environment by translating natural language commands into a specific navigation path. For example, R2R [mattersim, Matterport3D], RxR [rxr], REVERIE [REVERIE] and TOUCHDOWN [TOUCHDOWN] are benchmarks for this task.
Dialog-based Navigation. Dialog-based Navigation allows an agent to interact with a user and its physical world simultaneously, reasoning about ambiguous instructions through conversation to successfully navigate its environment. Related to it are RobotSlang [RobotSlang], R2H [R2H], CVDN [CVDN], UNMuTe [UNMuTe] and so on.
Evolution of Navigation Methods. Navigation tasks have seen a significant methodological evolution, moving from early, specialized systems to models powered by large language models (LLMs). Early non-LLM-based methods were typically trained on domain-specific data to create a direct mapping from sensory inputs to navigation actions [zhang2024visionandlanguagenavigationtodaytomorrow]. This reliance on implicit representations often led to models that learned spurious correlations between inputs and outputs, limiting their generalization to unseen environments and making their decisions difficult to interpret [zhang2025researchnavigationmethodsbased].
The advent of LLMs and VLMs has introduced a new paradigm. Pre-trained on vast, diverse datasets, these models bring a wealth of real-world commonsense and physical knowledge that dramatically improves navigation precision. This shift has enabled explicit reasoning, such as CoT [CoT], which not only enhances decision accuracy but also provides greater interpretability [zhang2025researchnavigationmethodsbased]. This has led to a new wave of methods which are shown in Table IV and Figure 8.
Figure 8: Overview of NavCoT (Figure used courtesy of [NavCoT]).
TABLE IV: Representative methods for navigation tasks: Non-LLM vs. LLM-based
@@
ObjectNav GOSE [chaplot2020objectgoalnavigationusing] RIM [RIM] LOAT [LOTA] LGR [LGR] CL-CoTNav [CL-CoTNav] ASCENT [ASCENT]
VLN BabyWalk [BabyWalk] ADAPT [ADAPT] HAMT [HAMT] EnvDrop [EnvDrop] DUET [DUET] NavGPT [NavGPT] NavCoT [NavCoT] VELMA [VELMA] NaVILA [NaVILA]
Dialog-based CMN [CMN] GVDN [GVDN] LLM- ContextBridge [ContextBridge] FLAME [FLAME]
However, these LLM-based approaches are not without their issues. They are susceptible to hallucinations, where incorrect physical reasoning can lead to failures, such as imagining a non-existent path through an obstacle [deng2025llmgoodpathplanner]. This fragility is further demonstrated by research showing that minor changes to a prompt can reduce a robot’s task success rate by nearly 20% [wu2025uncoveringfragilitytrustworthyllms]. A fundamental limitation is the representational bottleneck caused by converting rich physical world data into a token format, which can result in a significant loss of detail and the inability to reason about low-level preconditions for actions [RIM, chaplot2020objectgoalnavigationusing]. This is addressed by integrating world models, which can mentally simulate action outcomes, providing a crucial, grounded, and predictive layer that LLMs inherently lack [WebAgents]. This paradigm is also essential for bridging the "sim-to-real" gap, as a world model can learn a robust predictive representation of dynamics in simulation and transfer this knowledge to a real-world agent, as demonstrated by the TWIST [TWIST] framework’s teacher-student distillation approach.
@@ F.3 Autonomous Driving
Autonomous driving refers to a system’s capability to perform part or all of the dynamic driving task (DDT) on a sustained basis without direct human intervention. The DDT includes all of the real-time operational and tactical functions required to operate a vehicle, such as steering, acceleration, and braking. To achieve this, an autonomous vehicle’s closed-loop software system integrates a series of core functions: perception, prediction, planning, and control [thorn2018framework]. Perception involves collecting and processing real-time data from various sensors like cameras and LiDAR to detect objects and road conditions [Grigorescu_2019]. Prediction forecasts the behavior of other road users [Rudenko_2020]. Planning generates a safe and efficient path for the vehicle to follow. Control translates the planned path into physical commands for the vehicle’s actuators [paden2016surveymotionplanningcontrol]. These tasks are implemented through several architectural paradigms [bojarski2016endendlearningselfdriving]. Current autonomous driving technology methods mainly include: Rule-based Approaches, Learning-based Approaches, World Models and Generative Approaches, Hybrid Approaches, etc.
@@
Rule-based Approaches. Rule-based approaches rely on deterministic algorithms and explicit domain knowledge, encoding physics and traffic rules in interpretable models. Classical methods include the Intelligent Driver Model (IDM) [IDM] for car-following, RRT/RRT* [lavalle1998rapidly] for sampling-based trajectory generation, Model Predictive Control (MPC) [MPC] for optimization-based planning, and potential-field navigation [khatib1986real]. They are transparent and safe but struggle in dense and uncertain traffic. Benchmarks include CommonRoad [CommonRoad], nuPlan [nuplan], nuScenes [nuscenes], Waymo Open Dataset (WOD) [Waymo], CARLA [CARLA], and AirSim [AirSim].
@@
Learning-based Approaches. Learning-based approaches replace rules with neural policies trained from data. Imitation Learning (IL) follows expert trajectories (e.g., ChauffeurNet [ChauffeurNet]), Inverse Reinforcement Learning (IRL) recovers reward functions from demonstrations [Maximum], and Reinforcement Learning (RL) optimizes policies in simulators [kiran2021deepreinforcementlearningautonomous]. AlphaDrive [AlphaDrive] and AutoDrive-R2 [AutoDrive] extend RL-based training with physics-informed rewards, reinforcing the role of simulation and safety-oriented benchmarks in this category. More recent methods use transformer-based forecasting (e.g., LaneGCN [LaneGCN], Wayformer [Wayformer]).
@@
Generative Approaches. World models learn latent spatio-temporal representations of the driving environment, enabling rollouts of future states [World_models2018]. The Dreamer family (DriveDreamer [DriveDreamer] in Figure 9 and ReconDreamer [ReconDreamer]) capture compact latent dynamics for long-horizon simulation. Diffusion-based models include DrivingDiffusion [DrivingDiffusion], BEVControl [BEVControl], and Vista [Vista]. Transformer-based models such as GAIA-1/2 [GAIA-1] supports multimodal controllable scene generation. Occupancy-based models predict 4D voxelized space, such as OccWorld [OccWorld], RenderWorld [RenderWorld] and DOME [DOME] combine occupancy forecasting with planning via diffusion. Point-cloud generators like LiDARGen [LiDARGen] and the unified HERMES [HERMES] framework produce future LiDAR sweeps or merged 3D representations. Benchmarks in this domain include prediction and occupancy forecasting benchmarks (Occ3D-nuScenes/Waymo [Occ3D], OpenScene [OpenScene], Precog [PRECOG]), simulation benchmarks (CarlaSC [MotionSC], Bench2Drive [Bench2Drive]), and end-to-end planning benchmarks (nuPlan [nuplan], S2R-Bench [S2R-Bench]). These collectively test world models across scene generation, motion forecasting, and driving policy integration.
Figure 9: Overall framework of DriveDreamer (Figure used courtesy of [DriveDreamer]).
@@
Hybrid Approaches. Hybrid approaches combine rule-based reliability with learning-based adaptability and cognitive reasoning. Neural waypoint predictors refined by MPC or rule-based controllers augmented with learned forecasting exemplify this paradigm [Gupta_2023]. Cognitive reasoning extensions include DriveCoT [DriveCoT], which integrates Chain-of-Thought supervision into driving tasks, and PRIMEDrive-CoT [PRIMEDrive-CoT], which introduces uncertainty-aware reasoning. PlanAgent separates high-level planning from execution, while LeapVAD [LeapVAD] adopts a dual-process architecture inspired by human cognition. DriveLMM-o1 [DriveLMM-o1] and Reason2Drive [Reason2Drive] further contribute cognition-augmented datasets for training and evaluating reasoning.
@@
The current challenges in autonomous driving are fundamentally tied to the "long tail" problem of rare, high-impact edge cases that are difficult to encounter and address through traditional real-world data collection [wang2025terasimworldworldwidesafetycriticaldata, ren2025cosmosdrivedreamsscalablesyntheticdriving]. World models solve this by creating a learned, internal simulation of the environment, which functions as a "computational snow globe" where the AI can "mentally rehearse" actions [Think2Drive, hafner2024masteringdiversedomainsworld]. By generating and training in an unbounded number of synthetic, safety-critical scenarios, these models allow the autonomous system to anticipate and mitigate hazardous situations before they unfold in the physical world, ultimately shifting the paradigm from reactive to predictive autonomy and improving real-world safety [ding2023surveysafetycriticaldrivingscenario].
@@ H Conclusion
This survey presents a task-oriented taxonomy tracing physical AI development through four progressive capabilities: physical perception, physics reasoning, world modeling, and embodied interaction. Each stage builds upon the previous, advancing from passive observation to active physical comprehension. Despite progress in individual domains, most systems still rely on statistical pattern recognition rather than genuine physical understanding. Moving forward, the most promising path lies in hybrid approaches that integrate physics-grounded architectures, physics-informed training, and symbolic reasoning into unified frameworks. Such frameworks that traverse the entire progression from perception to interaction will enable AI systems to genuinely perceive, reason about, model, and interact with physical reality for safe and reliable deployment.